Find the percentage change in price in each of the following examples using the mid-point method. Instructions: Round your answers to two decimal places and include a negative sign where appropriate. a. The price of a $4 sandwich increases to $5 b. A sale discounts the price of a sofa from $750 to $500Suppose that when the average family income rises from $30,000 per year to $40,000 per year, the average familys purchases of toilet paper rise from 100 rolls to 105 rolls per year. Instructions: Round your answer to two decimal places and include a negative sign if appropriate. a. The income-elasticity of demand for toilet paper is 0.17 b. Toilet paper is a normal good D. loilet paper s i a normal good C. The demand for toilet paper is Uncome-inelasticThe factors that determine the price elasticity of demand include: Instructions: You may select more than one answer. Click the box with a check mark for correct answers and cli number of buyers. tastes and preferences. relative need and relative cost. time needed to adjust to price changes. availability of substitutes. technologyThe price elasticity of supply is always (Click to select) $) because the quantity supplied moves in the (Click to select) direction as the price. The price elasticity of demand is always (Click to select) because the quantity demanded always moves in the (Click to select) direction from the price.
Answer #1

a)The price of $4 sandwich increases to $5

Percentage change in price =P2-P1/(P2+P1)/2*100

PERCENTAGE CHANGE IN PRICE = 5-4/(5+4)/2*100

PERCENTAGE CHANGE IN PRICE =22.22%

b)A sale discounts price of a sofa from $750 to $ 500

PERCENTAGE CHANGE IN PRICE = P2-P1/(P2+P1)/2*100

PERCENTAGE CHANGE IN PRICE = 500-750/(500+750)/2*100

PERCENTAGE CHANGE IN PRICE =-250/625*100

PERCENTAGE CHANGE IN PRICE = -40.00%

INCOME ELASTICITY OF DEMAND FOR TOILET PAPER WHEN INCOME INCREASES FROM $30,000 TO $40,000,QUANTITY RISES FROM 100 TO 105 PER YEAR

a)income elasticity of toilet paper =% change in quantity demanded/% change in consumer income

INCOME ELASTICITY OF DEMAND (Ei) = Qf-Qi/(Qf+Qi/2)*100/If-Ii/(If+Ii/2)*100

(Ei)= 40000-30000/(40000+30000/2)*100/ 105-100(105+100/2)*100

(Ei)=0.17%

b)Toilet paper is a normal good.

(this is because,its income elasticity is positive)

c) Toilet paper is an income inelastic good

(this is because its income elasticity is greater than zero but it is less than one)

THE FACTORS THAT AFFECT THE PRICE ELASTICITY OF DEMAND INCLUDE

availability of substitutes

(when a commodity has large substitutes then a small rise in its price,induces consumers to buy its substitutes)

time needed to adjust to price changes

(consumers find it difficult to change their habits in a short period of time,in order to respond to the change in price time taken is thus an important factor)

tastes and preferences

(when the consumers prefer one good over the other or they are in the habit of using a certain good then change in price will not affect its demand)

relative need and relative cost

(when the need of the product is that of a luxury then the demand is price elastic,whereas necessities have an inelastic demand as well as relative cost of the commodity to the consumer,any product taking up a large portion of the income have elastic demand)

THE PRICE ELASTICITY OF SUPPLY IS ALWAYS POSITIVE BECAUSE THE QUANTITY SUPPLIED MOVES IN THE SAME DIRECTION AS THE PRICE

(this is because the law of supply states that supply of a commodity increases when there is an increase in its price other things remaining constant so there is a direct relationship between supply and price)

THE PRICE ELASTICITY OF DEMAND IS ALWAYS NEGATIVE BECAUSE THE QUANTITY DEMANDED MOVES IN THE OPPOSITE DIRECTION AS THE PRICE

(this is because the law of demand states that the quantity of a commodity demanded increases when its price decreases other things remaining constant so there is an inverse relationship between price and demand)